OpenStudy (anonymous):

Use a(t) = -9.8 meters per second per second as the acceleration due to gravity. (Neglect air resistance.) A baseball is thrown upward from a height of 3 meters with an initial velocity of 9 meters per second. Determine its maximum height. (Round your answer to two decimal places.)

OpenStudy (anonymous):

let a(t) = g= 9.8m/s^2 X1= 3m viy = 9m/s vfy = 0m/s at max height vf^2y = Vf^2y + 2gXo Xo = vf^2 -vi^2/2g Xo = 0-(9m/s)^2/2(-9.8m/s^2) Xo = 4.13m.. add the X1 (+3) = 7.13m I think that's the maximum height...